JEF meeting, September 20, 2024
From GlueXWiki
Meeting time and location:
- 9:30 am EST (JLab time) remote only
- Connection information
Agenda
Preparation of parts for ECAL -- Sasha
- - HV long cable: Sammy and Mindy are finishing up with assembling cables, 2 cables are left
- - HV distribution boards [1][2]
- - 5 (12 x 4) and 5 (6 x 8) boards have been received. Chris checked connections. All remaining PCB boards are expected to be in the lab on Sep 12.
- - Chris is finalizing design of the distribution boxes (each box will comprise 9 PCBs). He expects to submit PR during next week.
- - Fix front panel of two LV distribution boards [3]
- - new plates for the LV distribution boards should arrive today. Need to re-install 2 distribution boards and connect cables
Installation of the FCAL2 detector & integration to GlueX -- Sasha / Vlad / Hakob / Laveen
- Installation and testing dividers has been completed [4]
- - need to arrange cables on chains
- - need to arrange cables on cable trays - done [5]
- - preparing spare cables (7 short HV, 8 signal and LV)
- Installation of the light distribution board of the ECAL LMS [6]
- Implementing ECAL into the GlueX DAQ [7]
- Detector installation page
- Safety briefings
- Remaining installation tasks
- - Installing HV distribution boards (start switching the detector ON)
- - Installing dark room ceiling and insulation inside the room
- - Checking cooling system
Students -- Sasha
- - Expect one student from Wuhan, waiting for the US visa
- Restoring PS test setup
ECAL commissioning -- Sasha, All
Minutes
Attendees: Igal, Laveen, Phil, Liping, Vlad, Sasha, Zisis, Stjepan, Ilya, S.T.
- Waiting for cover panels for two LV boards
- No progress in HV distribution board design
- Cables organized on chains and in cable trays
- Installed final version of light distribution board for LMS on integrating sphere
- Looked at signals on some PMTS, pulses look good, but amplitudes somewhat small
- Need to increase light 10x
- Sasha implemented DAQ configuration for ECAL crates
- Optical links work, need to integrate LMS trigger
- Found faulty controller for ROCECAL3
- Dark room installation needs to be completed
- Checking of cooling system still needs to be completed before we can turn on the whole detector
- PS setup in the hall in the process of being restored (we can use this to study some small-scale prototypes)
- Started commissioning run plan